OpenStudy (anonymous):

I need help solving 4y-7=3y (what is y)

OpenStudy (uri):

4y-3y=7 y=7

OpenStudy (anonymous):

4y-7 = 3y subtract 3y to the left side and add 7 to the right side. You have 1y=7 divide7 by "1" y=7

OpenStudy (anonymous):

Osuokc im sorry could u explain it better

OpenStudy (anonymous):

Also I know I how to solve it once you tell me what y is but I just don't know how u got the answer

OpenStudy (uri):

Bring all the y values at one side,after equal to sign (=),when you shift it sign changes. So,4y-7 = 3y 4y-3y=7 y=7

OpenStudy (anonymous):

Ok I got it. Can you write a word problem that can be solve by this equation